Calculus II, Spring 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 06 and Week 07

1. Section 10.8: Problems 22, 23, 29, 31, 35, 37.

2. Section 11.7: Problems 3, 7, 9, 27, 29, 31, 35, 43.

3. Section 13.2: Problems 27, 29, 31, 34, 35, 45, 47.

4. Section 13.3: Problems 27, 35, 39.